OpenStudy (anonymous):

a^2 + 5a + 4 / a^3 * a^2 + 3a + 2 / a^2 - 2a

jimthompson5910 (jim_thompson5910):

oh ok, thx for clarifying

OpenStudy (anonymous):

np

jimthompson5910 (jim_thompson5910):

So the entire problem is \[\Large \left(\frac{a^2 + 5a + 4}{a^3} \right )\left(\frac{a^2 + 3a + 2}{a^2 - 2a} \right )\] correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

The first step is to factor a^2 + 5a + 4 a^2 + 5a + 4 a^2 + 4a + a + 4 (a^2+4a)+(a+4) a(a+4)+1(a+4) (a+1)(a+4) So a^2 + 5a + 4 factors to (a+1)(a+4)

jimthompson5910 (jim_thompson5910):

What do you get when you factor a^2 + 3a + 2 ??

OpenStudy (anonymous):

umm idk

jimthompson5910 (jim_thompson5910):

a^2 + 3a + 2 a^2 + 2a + a + 2 (a^2 + 2a) + (a + 2) a(a+2)+1(a+2) (a+1)(a+2) So a^2 + 3a + 2 factors to (a+1)(a+2)

jimthompson5910 (jim_thompson5910):

Then factor a^2 - 2a to get a(a-2)

OpenStudy (anonymous):

so the answer is a+4/a+2?

jimthompson5910 (jim_thompson5910):

Not quite

jimthompson5910 (jim_thompson5910):

wait a second, are we dividing these two big fractions or multiplying them?

OpenStudy (anonymous):

multiplying

jimthompson5910 (jim_thompson5910):

alright, one sec

OpenStudy (anonymous):

ok